# A Metric Hybrid Planning Approach to Solving Pandemic Planning Problems with Simple SIR Models

### v1

@@ Introduction
The brief description of the mathematical optimization model is as follows. The objective function (1) maximizes the total reward accumulated over the horizon. Constraint (2) sets the initial values of all state variables. Constraint (3) sets the values of state variables in the next step given the values of state and action variables in the current step. Constraints (4-6) enforce the instantanous, temporal and goal constraints. Constraints (7-9) set the domains of state, action and duration variables, respectively. Note that in the mathematical optimization model presented above, constraint (5) is an interval constraint that holds for all values of the interval $c^{t}\in[0,1]$. SCIPPlan uses a constraint generation framework that iteratively identifies unique violated values of $c^{t}\in[0,1]$ and generates constraint (5) with the indentified violated values of $c^{t}$. It has been shown that SCIPPlan terminates either with a solution to $\Pi$ with at most some constant $\gamma$ violation of constraint (5) or proves the infeasibility of $\Pi$ in finite number of constraint generation iterations. [^1] The proof assumes the temporal constraint function $C^{T}$ to be Lipschitz continuous. Other control functions can be used to derive domain-specific bounds based on the analysis of $C^{T}$, as we will show in section Theoretical Results. Next, we will describe the SIR model that will be used to construct the state transition function $T$ of our planning problem $\Pi$.
@@
Given the components of $\Pi$ are defined, SCIPPlan can compile $\Pi$ into a mathematical optimization model and solve it via constraint generation (i.e., previously described in section Solving the Metric Hybrid Planning Problem) However, the computational effectiveness of the underlying spatial branch-and-bound solver relies on its ability to effectively derive valid upper and lower bounds on the (nonlinear) expressions that make up the mathematical optimization model (i.e., the objective function (1) and the constraints (2-6)). Therefore, we will next derive valid inequalities in order to improve the computational effectiveness of SCIPPlan.
@@
where $e_{1}$, $e_{2}$ and $e_{3}$ denote the expressions $\frac{b(a_{1})}{b(a_{1})-c}$, $(b(a_{1})-c)\Delta^{t}$ and $\frac{c}{b(a_{1})-c}$, respectively. The brief description of the resulting mathematical optimization model is as follows. The objective function (17) minimizes the total number of lockdowns used over the horizon. Constraints (18-20) set the initial values of all state variables. Constraints (21-23) represent the state transition function $T$. Constraints (24-26) are the valid inequalities that are introduced in section Valid Inequalities. Constraint (27) is the temporal constraint. Constraints (28-29) are the goal state constraints. Finally, constraints (30-34) specify the domains of state and action variables, and the fixed duration of each step.
@@
where $\delta_{LB}$ and $\delta_{UB}$ denote the lower and upper bounds on the duration of each step, respectively.
@@
In order to solve Problem 1 with fixed step duration, we ran SCIPPlan with the value of $H$ such that $H\Delta$ is equal to 26 weeks. In order to solve Problem 2 with fixed step duration, we ran SCIPPlan with increasing values of horizon $H$ until either a solution is found or $H\Delta$ exceeded 52 weeks. If SCIPPlan returned a solution with $H\Delta$ that is less than 52 weeks, we verified that the temporal constraint $C^{T}(s^{t}_{1},s^{t}_{2},s^{t}_{3},a^{t}_{1},\Delta^{t})\leq 0$ holds for the remaining duration of the year without using additional actions (i.e., $a_{1}^{t}=0$), otherwise we ran SCIPPlan with the incremented value of $H$. Finally, in order to solve Problem 1 with variable step duration, we have used $\delta_{LB}=7$ and $\delta_{UB}=28$ as the lower and upper bounds of the duration of each step with horizon $H=8$.
@@
We begin our analysis with the evaluation of the effect of valid inequalities that are introduced in section Valid Inequalities (i.e., constraints (24-26)) on the overall runtime performance of SCIPPlan. Figure 2 visualizes the logarithmic runtime comparison between using SCIPPlan with and without the valid inequalities over all instances. In figure 2, each data point represents an instance that corresponds to a unique combination of the parameters provided in Table 1. The inspection of figure 2 clearly highlights the benefit of including the valid inequalities as a part of the planning problems. On average, we observe that the addition of the valid inequalities improves the runtime performance of SCIPlan by around one to two orders of magnitude. As a result, we report results on the version of the planning problems with the valid inequalities in the remaining of this section.
@@
In section Theoretical Results, we have shown theoretical results on the finiteness and solution quality of SCIPPlan for solving the pandemic planning problem. Our theoretical results yield strong guarantees on the solutions provided by SCIPPlan such that the temporal constraint is satisfied with relatively small values of $\gamma$ where $\gamma$ is 1.00 for $K=200$ and $\gamma$ is $1.25$ for $K=250$.
In section Experimental Results, we have shown the computational performance of our approach to control a pandemic and other related systems. Our experimental results demonstrate the computational benefit of deriving valid inequalities for the planning problem. Some of these valid inequalities (e.g., monotonicity of state variables) are common in many other continuous time decision making problems where the state evolution is governed by a system of partial differential equations. For future work, we plan to investigate the automation of the process of valid inequality derivation for similar metric hybrid planning problems. Further, our experimental results also demonstrate the computational viability of using the exact solution equations of the SIR model (i.e., equations (10- 12)) to construct a nonlinear state transition function $T$ using both exponential and logarithmic expressions, and perform formal reasoning over continuous time. Our experimental results compliment the success of the existing decision making systems that allow for continuous time decision making under additional restrictive assumptions (e.g., the assumption that the state transition function $T$ is piecewise linear (Shin and Davis 2005; Coles et al. 2012; Chen, Williams, and Fan 2021), polynomial (Cashmore et al. 2016) etc.). Finally, we highlight the importance of our approach given the availability of machine learning techniques (Raissi, Perdikaris, and Karniadakis 2019; Karniadakis et al. 2021) that can successfully approximate the solution equations of many systems of partial differential equations using data. For future work, we plan to apply our approach to control a pandemic and other related systems using SCIPPlan based on learned models (Wu, Say, and Sanner 2017; Say and Sanner 2018b; Say, Sanner, and Thiébaux 2019; Say and Sanner 2020; Say et al. 2020; Say 2020; Wu, Say, and Sanner 2020; Say 2021).
@@ Conclusion
Ethics Statement
This work studies the computational effectiveness of using the solution equations of a SIR model, and does not constitute as health advice and does not take the place of consulting with the experts of the relevant fields.

### v2

@@ Introduction
The brief description of the mathematical optimization model is as follows. The objective function (1) maximizes the total reward accumulated over the horizon. Constraint (2) sets the initial values of all state variables. Constraint (3) sets the values of state variables in the next step given the values of state and action variables in the current step. Constraints (4-6) enforce the instantanous, temporal and goal constraints. Constraints (7-9) set the domains of state, action and duration variables, respectively. Note that in the mathematical optimization model presented above, constraint (5) is an interval constraint that holds for all values of the interval $c^{t}\in[0,1]$. SCIPPlan uses a constraint generation framework that iteratively identifies unique violated values of $c^{t}\in[0,1]$ and generates constraint (5) with the indentified violated values of $c^{t}$. It has been shown that SCIPPlan terminates either with a solution to $\Pi$ with at most some constant $\gamma$ violation of constraint (5) or proves the infeasibility of $\Pi$ in finite number of constraint generation iterations. [^1] The proof assumes the temporal constraint function $C^{T}$ to be Lipschitz continuous. Other control functions can be used to derive domain-specific bounds based on the analysis of $C^{T}$, as we will show in the Theoretical Results section. Next, we will describe the SIR model that will be used to construct the state transition function $T$ of our planning problem $\Pi$.
@@
Given the components of $\Pi$ are defined, SCIPPlan can compile $\Pi$ into a mathematical optimization model and solve it via constraint generation (i.e., previously described in the Solving the Metric Hybrid Planning Problem section). However, the computational effectiveness of the underlying spatial branch-and-bound solver relies on its ability to effectively derive valid upper and lower bounds on the (nonlinear) expressions that make up the mathematical optimization model (i.e., the objective function (1) and the constraints (2-6)). Therefore, we will next derive valid inequalities in order to improve the computational effectiveness of SCIPPlan.
@@
where $e_{1}$, $e_{2}$ and $e_{3}$ denote the expressions $\frac{b(a_{1})}{b(a_{1})-c}$, $(b(a_{1})-c)\Delta^{t}$ and $\frac{c}{b(a_{1})-c}$, respectively. The brief description of the resulting mathematical optimization model is as follows. The objective function (17) minimizes the total number of lockdowns used over the horizon. Constraints (18-20) set the initial values of all state variables. Constraints (21-23) represent the state transition function $T$. Constraints (24-26) are the valid inequalities that are introduced in the Valid Inequalities section. Constraint (27) is the temporal constraint. Constraints (28-29) are the goal state constraints. Finally, constraints (30-34) specify the domains of state and action variables, and the fixed duration of each step.
@@
where $\delta_{LB}$ and $\delta_{UB}$ denote the lower and upper bounds on the duration of each step, respectively. Finally, in order to cover a fixed total duration, we add the following constraint:
$\sum_{t=1}^{H}\Delta^{t}=F$ (40)
where $F$ denotes the fixed total duration.
@@
In order to solve Problem 1 with fixed step duration, we ran SCIPPlan with the value of $H$ such that $H\Delta$ is equal to 24 weeks. In order to solve Problem 2 with fixed step duration, we ran SCIPPlan with increasing values of horizon $H$ until either a solution is found or $H\Delta$ exceeded 52 weeks. If SCIPPlan returned a solution with $H\Delta$ that is less than 52 weeks, we verified that the temporal constraint $C^{T}(s^{t}_{1},s^{t}_{2},s^{t}_{3},a^{t}_{1},\Delta^{t})\leq 0$ holds for the remaining duration of the year without using additional actions (i.e., $a_{1}^{t}=0$), otherwise we ran SCIPPlan with the incremented value of $H$. Finally, in order to solve Problem 1 with variable step duration, we have used $\delta_{LB}=7$ and $\delta_{UB}=28$ as the lower and upper bounds of the duration of each step with horizon $H=8$ and fixed total duration $F=168$ (i.e., 24 weeks).
@@
We begin our analysis with the evaluation of the effect of valid inequalities that are introduced in the Valid Inequalities section (i.e., constraints (24-26)) on the overall runtime performance of SCIPPlan. Figure 2 visualizes the logarithmic runtime comparison between using SCIPPlan with and without the valid inequalities over all instances. In figure 2, each data point represents an instance that corresponds to a unique combination of the parameters provided in Table 1. The inspection of figure 2 clearly highlights the benefit of including the valid inequalities as a part of the planning problems. On average, we observe that the addition of the valid inequalities improves the runtime performance of SCIPlan by around one to two orders of magnitude. As a result, we report results on the version of the planning problems with the valid inequalities in the remaining of this section.
@@
In the Theoretical Results section, we have shown theoretical results on the finiteness and solution quality of SCIPPlan for solving the pandemic planning problem. Our theoretical results yield strong guarantees on the solutions provided by SCIPPlan such that the temporal constraint is satisfied with relatively small values of $\gamma$ where $\gamma$ is 1.00 for $K=200$ and $\gamma$ is $1.25$ for $K=250$.
In the Experimental Results section, we have shown the computational performance of our approach to control a pandemic and other related systems. Our experimental results demonstrate the computational benefit of deriving valid inequalities for the planning problem. Some of these valid inequalities (e.g., monotonicity of state variables) are common in many other continuous time decision making problems where the state evolution is governed by a system of partial differential equations. For future work, we plan to investigate the automation of the process of valid inequality derivation for similar metric hybrid planning problems. Further, our experimental results also demonstrate the computational viability of using the exact solution equations of the SIR model (i.e., equations (10- 12)) to construct a nonlinear state transition function $T$ using both exponential and logarithmic expressions, and perform formal reasoning over continuous time. Our experimental results compliment the success of the existing decision making systems that allow for continuous time decision making under additional restrictive assumptions (e.g., the assumption that the state transition function $T$ is piecewise linear (Shin and Davis 2005; Coles et al. 2012; Chen, Williams, and Fan 2021), polynomial (Cashmore et al. 2016) etc.). Finally, we highlight the importance of our approach given the availability of machine learning techniques (Raissi, Perdikaris, and Karniadakis 2019; Karniadakis et al. 2021) that can successfully approximate the solution equations of many systems of partial differential equations using data. For future work, we plan to apply our approach to control a pandemic and other related systems using SCIPPlan based on learned models (Wu, Say, and Sanner 2017; Say and Sanner 2018b; Say, Sanner, and Thiébaux 2019; Say and Sanner 2020; Say et al. 2020; Say 2020; Wu, Say, and Sanner 2020; Say 2021).
@@ Conclusion
Disclaimer
This work studies the computational effectiveness of using the solution equations of an SIR model, and does not constitute as health advice and does not take the place of consulting with the experts of the relevant fields.